# Revisiting Randomization in Greedy Model Search

### v1

@@ A.2 Proofs for Asymptotic RGS Weights
$\frac{1}{1+e^{-\alpha d}}\leq\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1)}}.$ (31)
Additionally, using the bounds from (31), we have
$\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1)}}\leq\check{w}_{d+1}^{\gamma},$ (32)
demonstrating that $\check{w}_{d}^{\gamma}$ is non-decreasing for $d\in\mathbb{Z}$. The inequality $\check{w}_{d}^{\gamma}\geq\frac{1}{1+e^{-\alpha(d+1/2)}}$ for $d\geq 0$ follows directly from (38). Furthermore, we obtain $\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1/2)}}$ for $d<0$ by using the relation $\check{w}_{d}^{\gamma}=1-\check{w}_{-(d+1)}^{\gamma}$.
To prove property (d), note that property (i) in the footnote follows immediately from (32). Next, we prove properties (ii) and (iii) in the footnote. To this end, using the two different series representations (15) and (26) of $\check{w}_{d}^{\gamma}$ for integer $d$, we extend the definition of $\check{w}_{d}^{\gamma}$ to all real $d$ as follows. When $d<-1/2$, let $\check{w}_{d}^{\gamma}=\sum_{i=1}^{\infty}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}$ and when $d\geq-1/2$, let $\check{w}_{d}^{\gamma}=\sum_{i=0}^{\infty}(-1)^{i}e^{-\alpha i(d+i/2+1/2)}$. For $d>-1/2$, note that $\check{w}_{d}^{\gamma}$ has second derivative
$\frac{\partial^{2}}{\partial d^{2}}\check{w}_{d}^{\gamma}=\alpha^{2}\sum_{i=0}^{\infty}(-1)^{i}i^{2}e^{-\alpha i(d+i/2+1/2)}=\alpha^{2}\sum_{i=0}^{\infty}(-1)^{i}z^{2i}i^{2}q^{i^{2}},$ (33)
and $z,q\in(0,1)$, where $q=e^{-\alpha/2}$ and $z=e^{-(\alpha/2)(d+1/2)}$. Now, consider the following identity from [17, Proposition 1.1]:
$\sum_{i=0}^{\infty}(-1)^{i}z^{2i}q^{i^{2}}=\frac{1}{1+z}\sum_{i=0}^{\infty}z^{i}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}},\quad|z|<1,\;|q|\leq 1.$ (34)
Differentiating both sides of (34) with respect to $q$ gives
$\sum_{i=0}^{\infty}(-1)^{i}z^{2i}i^{2}q^{i^{2}-1}=\frac{1}{1+z}\sum_{i=1}^{\infty}z^{i}\frac{\partial}{\partial q}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}}.$ (35)
By the product rule and quotient rule on (35),
$\frac{\partial}{\partial q}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}}=\sum_{l=0}^{i-1}\frac{q^{l-1}z(q^{l+1}z-l(q+1)-q)}{(1+zq^{l+1})^{2}}\prod_{0\leq j\leq i-1,\,j\neq l}\frac{1-zq^{j}}{1+zq^{j+1}}.$ (36)
Since $q^{l+1}z-l(q+1)-q\leq 0$ for all $l\geq 0$ and $(1-zq^{j})/(1+zq^{j+1})\geq 0$ for all $j\geq 0$ when $z,q\in[0,1)$, it follows that (36) is non-positive, implying the same for (35). Returning to (33), we deduce that $\check{w}_{d}^{\gamma}$ is concave for all $d>-1/2$. An analogous argument shows that $\check{w}_{d}^{\gamma}$ is convex for $d<-1/2$. This proves properties (ii) and (iii) in the footnote when $d\neq-1,0$. For the $d=-1,0$ cases, we have from the just established convexity/concavity and continuity that
$\frac{1}{3}\check{w}_{-2}^{\gamma}+\frac{2}{3}\lim_{d\uparrow-1/2}\check{w}_{d}^{\gamma}$ $\geq\check{w}_{-1}^{\gamma},$ (37)
@@
where $\lim_{d\uparrow-1/2}\check{w}_{d}^{\gamma}=1-\check{w}_{-1/2}^{\gamma}$. We will show that these equations (37) together imply properties (ii) and (iii) in the footnote when $d=-1,0$. Expanding the power series (34), ones obtains
$\check{w}_{d}^{\gamma}=\frac{1}{1+z}\sum_{i=0}^{\infty}z^{i}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}}\geq\frac{1}{1+z}\sum_{i=0}^{\infty}z^{i}\prod_{j=0}^{i-1}\frac{1-z}{1+z}=\frac{1}{1+z^{2}}=\frac{1}{1+e^{-\alpha(d+1/2)}}.$ (38)
Thus, we have $\check{w}_{-1/2}^{\gamma}\geq 1/2$ and hence $\check{w}_{-1/2}^{\gamma}\geq\lim_{d\uparrow-1/2}\check{w}_{d}^{\gamma}$. Thus, from (37), we obtain
$3\check{w}_{-1}^{\gamma}+\check{w}_{1}^{\gamma}\leq 3\check{w}_{0}^{\gamma}+\check{w}_{-2}^{\gamma}.$ (39)
Finally, the identity $\check{w}_{d}^{\gamma}=1-\check{w}_{-(d+1)}^{\gamma}$ for $d\in\mathbb{Z}$ from part (c) of Theorem 4.2 applied to (39) yields properties (ii) and (iii) in the footnote when $d=-1,0$. ∎
Proof of Theorem 4.3.
We leave the proof to the end of Section A.3. ∎
A.3 Proof for Error Improvement
Proof of Theorem 3.2.
@@ A.3 Proof for Error Improvement
$\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}=\sum_{j=k+1}^{p}\hat{\beta}^{2}_{j,p}\big{(}1-\big{(}-w_{j}^{k,m,p}\big{)}^{2}\big{)}-\sum_{j=1}^{k}\hat{\beta}^{2}_{j,p}\big{(}1-w_{j}^{k,m,p}\big{)}^{2}.$ (40)
Let $m_{p}=\lfloor\gamma p\rfloor$, where $10<\gamma<$ 1. Then, using (40), we have
@@
$\qquad\geq\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big{(}1-\big{(}-w_{j}^{k,m_{p},p}\big{)}^{2}\big{)}-\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big{(}1-w_{j}^{k,m_{p},p}\big{)}^{2}.$ (41)
To simplify further, we use Chebyshev’s sum inequality [30], noting that $\mathbb{E}[\hat{\beta}^{2}_{j,p}]$ is non-increasing in $j$ by definition and $w_{j}^{k,m_{p},p}$ is non-increasing in $j$ by Theorem 4.1. Thus, the second term in (41) is upper bounded by
$\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big{(}1-w_{j}^{k,m_{p},p}\big{)}^{2}$ $\leq\Bigg{(}\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\Bigg{)}\Bigg{(}\sum_{j=1}^{k}\big{(}1-w_{j}^{k,m_{p},p}\big{)}^{2}\Bigg{)},$ (42)
and the first term in (41) is lower bounded by
@@
$\qquad\qquad\geq\Bigg{(}\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\Bigg{)}\Bigg{(}\sum_{j=k+1}^{2k}\big{(}1-\big{(}1-w_{j}^{k,m_{p},p}\big{)}^{2}\big{)}\Bigg{)}.$ (43)
@@
$\sum_{j=k+1}^{2k}\big{(}1-\big{(}-w_{j}^{k,m_{p},p}\big{)}^{2}\big{)}\geq(1-\epsilon)\sum_{j=k+1}^{2k}\big{(}1-\big{(}-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\big{)},$ (44)
@@
$\sum_{j=1}^{k}\big{(}1-w_{j}^{k,m_{p},p}\big{)}^{2}\leq(1+\epsilon)\sum_{j=1}^{k}\big{(}1-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}.$ (45)
For simplicity of notation, throughout the proof of this theorem, we use the notation “$A\lesssim B$” to mean that there exists a constant $C$ depending on $k$ such that $A\leq CB$. To prove the claim, we first show that when $p(1-\gamma_{p})^{2}$ is sufficiently large,
$\frac{\Big{|}\sum_{j=k+1}^{2k}\big{(}1-\big{(}-w_{j}^{k,m_{p},p}\big{)}^{2}\big{)}-\sum_{j=k+1}^{2k}\big{(}1-\big{(}-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\big{)}\Big{|}}{\sum_{j=k+1}^{2k}\big{(}1-\big{(}-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\big{)}}\lesssim\frac{1}{p(1-\gamma_{p})^{2}}.$ (46)
If (46) holds, then it is clear that the claim for (44) holds.
To prove (46), we revisit the proof of Lemma A.1. Note that
@@
$\frac{1+p-m_{p}}{p-j-m_{p}+1}-1=\frac{j}{p-j-m_{p}+1}\lesssim\frac{1}{p(1-\gamma_{p})}.$ (47)
@@
$\Bigg{|}\sum_{j=k+1}^{2k}\big{(}1-\big{(}-w_{j}^{k,m_{p},p}\big{)}^{2}\big{)}-\sum_{j=k+1}^{2k}\big{(}1-\big{(}-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\big{)}\Bigg{|}\lesssim\frac{1}{p(1-\gamma_{p})}.$ (48)
@@
$\sum_{j=k+1}^{2k}\big{(}1-\big{(}-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\big{)}\gtrsim 1-\gamma_{p}.$ (49)
To lower bound $\sum_{j=k+1}^{2k}\big{(}1-\big{(}-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\big{)}$ from (49), note that
$1-\big{(}1-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}=\tilde{w}_{j}^{k,\gamma_{p}}\big{(}2-\tilde{w}_{j}^{k,\gamma_{p}}\big{)},$ (50)
@@
$\tilde{w}_{j}^{k,\gamma_{p}}\geq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha k}+e^{-\alpha_{p}(k-j)}}\geq e^{-\alpha_{p}(j-k)}\frac{1-e^{-\alpha_{p}k}}{1+e^{-\alpha_{p}}},\quad j\geq k+1.$ (51)
@@
$\tilde{w}_{j}^{k,\gamma_{p}}\leq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}k}+e^{-\alpha_{p}(k-j+1)}}\leq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}k}+1}\leq 1/2.$ (52)
From (51) and (52), we obtain
$1\displaystyle 1-(1-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}=\tilde{w}_{j}^{k,\gamma_{p}}(2-\tilde{w}_{j}^{k,\gamma_{p}})\geq\frac{3}{2}e^{-\alpha_{p}(j-k)}\frac{1-e^{-\alpha_{p}k}}{1+e^{-\alpha_{p}}},\quad j\geq k+$ 1 . (53)
Substituting in (53), we thus have
$\sum_{j=k+1}^{2k}\big{(}1-\big{(}-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\big{)}\geq\frac{3}{2}e^{-\alpha_{p}}\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}}}\frac{1-e^{-\alpha_{p}k}}{1+e^{-\alpha_{p}}}=\frac{3}{2}\frac{e^{-\alpha_{p}}(1-e^{-\alpha_{p}k})^{2}}{1-e^{-2\alpha_{p}}}.$ (54)
@@
which establishes (49). Combining (48) and (49) proves (46), which in turn confirms (44).
To prove (45), we show that when $\gamma_{p}\gtrsim 1$,
$\sum_{j=1}^{k}\big{(}1-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\gtrsim\big{(}1-\gamma_{p})^{2}.$ (55)
When $p(1-\gamma_{p})^{3}$ is sufficiently large, if (55) holds, then using the fact that $\big{|}w_{j}^{k,m_{p},p}-\tilde{w}_{j}^{k,\gamma_{p}}\big{|}\lesssim\frac{1}{p(1-\gamma_{p})}$, we know
@@
Noting that $\gamma_{p}\gtrsim$ implies $\alpha_{p}\gtrsim 1$, we obtain (55).
Therefore, the claim is proven for (44) and (45). To upper bound the sum $\sum_{j=1}^{k}\big{(}1-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}$ from (45), we use the following rearranged version of recurrence (24):
$1-\tilde{w}_{j}^{k,\gamma_{p}}=e^{-\alpha_{p}k}(1-\tilde{w}_{j-1}^{k,\gamma_{p}})+e^{-\alpha_{p}(k-j+1)}\tilde{w}_{j-1}^{k,\gamma_{p}}.$ (56)
Upon squaring both sides of (56), we obtain
@@
$\qquad 2e^{-\alpha_{p}(2k-j+1)}\tilde{w}_{j-1}^{k,\gamma_{p}}(1-\tilde{w}_{j-1}^{k,\gamma_{p}})+e^{-2\alpha_{p}(k-j+1)}\big{(}\tilde{w}_{j-1}^{k,\gamma_{p}}\big{)}^{2}.$ (57)
Summing (A.3) over $1\leq j\leq k$, using the Cauchy-Schwarz inequality, and the fact from Theorem 4.1 that $\tilde{w}_{j-1}^{k,\gamma_{p}}\geq\tilde{w}_{j}^{k,\gamma_{p}}$, we obtain
@@
$\qquad+2e^{-\alpha_{p}k}\sqrt{\sum_{j=1}^{k}\big{(}1-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}\big{(}\tilde{w}_{j-1}^{k,\gamma_{p}}\big{)}^{2}}$ (58)
@@
Solving the quadratic inequality (A.3) yields
$\sum_{j=1}^{k}\big{(}1-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\leq\frac{\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}\big{(}\tilde{w}_{j-1}^{k,\gamma_{p}}\big{)}^{2}}{(1-e^{-\alpha_{p}k})^{2}}.$ (59)
Now, from the upper bound (14), we have $\big{(}\tilde{w}_{j-1}^{k,\gamma_{p}}\big{)}^{2}\leq(1-e^{-\alpha_{p}k})^{2}$ for all $j\geq 1$. Substituting this into (59), we obtain the bound
$\sum_{j=1}^{k}\big{(}1-\tilde{w}_{j}^{k,\gamma_{p}}\big{)}^{2}\leq\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}=\frac{e^{-2\alpha_{p}}(1-e^{-2\alpha_{p}k})}{1-e^{-2\alpha_{p}}}.$ (60)
@@
$A=\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}],\qquad B=\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}],$ (61)
and note that $A\geq B$ by applying Definition 3.1. Using (61) and by the previous inequalities in (60) and (54), the training error gap (41) is at least
@@
$\qquad\geq\frac{3}{2}(1-\epsilon)Be^{-\alpha_{p}}\frac{1-e^{-\alpha_{0}}}{1+e^{-\alpha_{0}}}-(1+\epsilon)Ae^{-2\alpha_{p}},$ (62)

### v3

@@ Acknowledgements
$\frac{1}{1+e^{-\alpha d}}\leq\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1)}}.$ (S.11)
Additionally, using the bounds from (S.11), we have
$\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1)}}\leq\check{w}_{d+1}^{\gamma},$ (S.12)
demonstrating that $\check{w}_{d}^{\gamma}$ is non-decreasing for $d\in\mathbb{Z}$. The inequality $\check{w}_{d}^{\gamma}\geq\frac{1}{1+e^{-\alpha(d+1/2)}}$ for $d\geq 0$ follows directly from (S.18). Furthermore, we obtain $\check{w}_{d}^{\gamma}\leq\frac{1}{1+e^{-\alpha(d+1/2)}}$ for $d<0$ by using the relation $\check{w}_{d}^{\gamma}=1-\check{w}_{-(d+1)}^{\gamma}$.
To prove property (d), note that property (i) in the footnote follows immediately from (S.12). Next, we prove properties (ii) and (iii) in the footnote. To this end, using the two different series representations (11) and (S.6) of $\check{w}_{d}^{\gamma}$ for integer $d$, we extend the definition of $\check{w}_{d}^{\gamma}$ to all real $d$ as follows. When d<1/2d<-1/2, let $\check{w}_{d}^{\gamma}=\sum_{i=1}^{\infty}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}$ and when $d\geq-1/2$, let $\check{w}_{d}^{\gamma}=\sum_{i=0}^{\infty}(-1)^{i}e^{-\alpha i(d+i/2+1/2)}$. For d>1/2d>-1/2, note that $\check{w}_{d}^{\gamma}$ has second derivative
$\frac{\partial^{2}}{\partial d^{2}}\check{w}_{d}^{\gamma}=\alpha^{2}\sum_{i=0}^{\infty}(-1)^{i}i^{2}e^{-\alpha i(d+i/2+1/2)}=\alpha^{2}\sum_{i=0}^{\infty}(-1)^{i}z^{2i}i^{2}q^{i^{2}},$ (S.13)
and $z,q\in(0,1)$, where $q=e^{-\alpha/2}$ and $z=e^{-(\alpha/2)(d+1/2)}$. Now, consider the following identity from (coogan2003q, Proposition 1.1):
$\sum_{i=0}^{\infty}(-1)^{i}z^{2i}q^{i^{2}}=\frac{1}{1+z}\sum_{i=0}^{\infty}z^{i}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}},\quad|z|<1,\;|q|\leq 1.$ (S.14)
Differentiating both sides of (S.14) with respect to $q$ gives
$\sum_{i=0}^{\infty}(-1)^{i}z^{2i}i^{2}q^{i^{2}-1}=\frac{1}{1+z}\sum_{i=1}^{\infty}z^{i}\frac{\partial}{\partial q}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}}.$ (S.15)
By the product rule and quotient rule on (S.15),
$\frac{\partial}{\partial q}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}}=\sum_{l=0}^{i-1}\frac{q^{l-1}z(q^{l+1}z-l(q+1)-q)}{(1+zq^{l+1})^{2}}\prod_{0\leq j\leq i-1,\,j\neq l}\frac{1-zq^{j}}{1+zq^{j+1}}.$ (S.16)
Since $q^{l+1}z-l(q+1)-q\leq 0$ for all $l\geq 0$ and $(1-zq^{j})/(1+zq^{j+1})\geq 0$ for all $j\geq 0$ when $z,q\in[0,1)$, it follows that (S.16) is non-positive, implying the same for (S.15). Returning to (S.13), we deduce that $\check{w}_{d}^{\gamma}$ is concave for all d>1/2d>-1/2. An analogous argument shows that $\check{w}_{d}^{\gamma}$ is convex for d<1/2d<-1/2. This proves properties (ii) and (iii) in the footnote when $d\neq-1,0$. For the d=1,0d=-1,0 cases, we have from the just established convexity/concavity and continuity that
$\frac{1}{3}\check{w}_{-2}^{\gamma}+\frac{2}{3}\lim_{d\uparrow-1/2}\check{w}_{d}^{\gamma}$ $\geq\check{w}_{-1}^{\gamma},$ (S.17)
@@
where $\lim_{d\uparrow-1/2}\check{w}_{d}^{\gamma}=1-\check{w}_{-1/2}^{\gamma}$. We will show that these equations (S.17) together imply properties (ii) and (iii) in the footnote when d=1,0d=-1,0. Expanding the power series (S.14), ones obtains
$\check{w}_{d}^{\gamma}=\frac{1}{1+z}\sum_{i=0}^{\infty}z^{i}\prod_{j=0}^{i-1}\frac{1-zq^{j}}{1+zq^{j+1}}\geq\frac{1}{1+z}\sum_{i=0}^{\infty}z^{i}\prod_{j=0}^{i-1}\frac{1-z}{1+z}=\frac{1}{1+z^{2}}=\frac{1}{1+e^{-\alpha(d+1/2)}}.$ (S.18)
Thus, we have $\check{w}_{-1/2}^{\gamma}\geq 1/2$ and hence $\check{w}_{-1/2}^{\gamma}\geq\lim_{d\uparrow-1/2}\check{w}_{d}^{\gamma}$. Thus, from (S.17), we obtain
$3\check{w}_{-1}^{\gamma}+\check{w}_{1}^{\gamma}\leq 3\check{w}_{0}^{\gamma}+\check{w}_{-2}^{\gamma}.$ (S.19)
Finally, the identity $\check{w}_{d}^{\gamma}=1-\check{w}_{-(d+1)}^{\gamma}$ for $d\in\mathbb{Z}$ from part (c) of Theorem 3.2 applied to (S.19) yields properties (ii) and (iii) in the footnote when d=1,0d=-1,0. ∎
S.1.3 Proof of Theorem 4.1
@@
$\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}=\sum_{j=k+1}^{p}\hat{\beta}^{2}_{j,p}\big(1-\big(-w_{j}^{k,m,p}\big)^{2}\big)-\sum_{j=1}^{k}\hat{\beta}^{2}_{j,p}\big(1-w_{j}^{k,m,p}\big)^{2}.$ (S.20)
Let $m_{p}=\lfloor\gamma p\rfloor$, where $0<\gamma<1$. Then, using (S.20), we have
@@
$\qquad\geq\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big(1-\big(-w_{j}^{k,m_{p},p}\big)^{2}\big)-\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big(1-w_{j}^{k,m_{p},p}\big)^{2}.$ (S.21)
To simplify further, we use Chebyshev’s sum inequality (hardy1988inequalities), noting that $\mathbb{E}[\hat{\beta}^{2}_{j,p}]$ is non-increasing in $j$ by definition and $w_{j}^{k,m_{p},p}$ is non-increasing in $j$ by Theorem 3.1. Thus, the second term in (S.21) is upper bounded by
$\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\big(1-w_{j}^{k,m_{p},p}\big)^{2}$ $\leq\Bigg(\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\Bigg)\Bigg(\sum_{j=1}^{k}\big(1-w_{j}^{k,m_{p},p}\big)^{2}\Bigg),$ (S.22)
and the first term in (S.21) is lower bounded by
@@
$\qquad\qquad\geq\Bigg(\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]\Bigg)\Bigg(\sum_{j=k+1}^{2k}\big(1-\big(1-w_{j}^{k,m_{p},p}\big)^{2}\big)\Bigg).$ (S.23)
@@
$\sum_{j=k+1}^{2k}\big(1-\big(-w_{j}^{k,m_{p},p}\big)^{2}\big)\geq(1-\epsilon)\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big),$ (S.24)
@@
$\sum_{j=1}^{k}\big(1-w_{j}^{k,m_{p},p}\big)^{2}\leq(1+\epsilon)\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}.$ (S.25)
For simplicity of notation, throughout the proof of this theorem, we use the notation $A\lesssim B$” to mean that there exists a constant $C$ depending on $k$ such that $A\leq CB$. To prove the claim, we first show that when $p(1-\gamma_{p})^{2}$ is sufficiently large,
$\frac{\Big|\sum_{j=k+1}^{2k}\big(1-\big(-w_{j}^{k,m_{p},p}\big)^{2}\big)-\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)\Big|}{\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)}\lesssim\frac{1}{p(1-\gamma_{p})^{2}}.$ (S.26)
If (S.26) holds, then it is clear that the claim for (S.24) holds.
To prove (S.26), we revisit the proof of Lemma S.1.1. Note that
@@
$\frac{1+p-m_{p}}{p-j-m_{p}+1}-1=\frac{j}{p-j-m_{p}+1}\lesssim\frac{1}{p(1-\gamma_{p})}.$ (S.27)
@@
$\Bigg|\sum_{j=k+1}^{2k}\big(1-\big(-w_{j}^{k,m_{p},p}\big)^{2}\big)-\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)\Bigg|\lesssim\frac{1}{p(1-\gamma_{p})}.$ (S.28)
@@
$\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)\gtrsim 1-\gamma_{p}.$ (S.29)
To lower bound $\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)$ from (S.29), note that
$1-\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}=\tilde{w}_{j}^{k,\gamma_{p}}\big(2-\tilde{w}_{j}^{k,\gamma_{p}}\big),$ (S.30)
@@
$\tilde{w}_{j}^{k,\gamma_{p}}\geq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha k}+e^{-\alpha_{p}(k-j)}}\geq e^{-\alpha_{p}(j-k)}\frac{1-e^{-\alpha_{p}k}}{1+e^{-\alpha_{p}}},\quad j\geq k+1.$ (S.31)
@@
$\tilde{w}_{j}^{k,\gamma_{p}}\leq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}k}+e^{-\alpha_{p}(k-j+1)}}\leq\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}k}+1}\leq 1/2.$ (S.32)
From (S.31) and (S.32), we obtain
$1-(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}=\tilde{w}_{j}^{k,\gamma_{p}}(2-\tilde{w}_{j}^{k,\gamma_{p}})\geq\frac{3}{2}e^{-\alpha_{p}(j-k)}\frac{1-e^{-\alpha_{p}k}}{1+e^{-\alpha_{p}}},\quad j\geq k+1.$ (S.33)
Substituting in (S.33), we thus have
$\sum_{j=k+1}^{2k}\big(1-\big(-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\big)\geq\frac{3}{2}e^{-\alpha_{p}}\frac{1-e^{-\alpha_{p}k}}{1-e^{-\alpha_{p}}}\frac{1-e^{-\alpha_{p}k}}{1+e^{-\alpha_{p}}}=\frac{3}{2}\frac{e^{-\alpha_{p}}(1-e^{-\alpha_{p}k})^{2}}{1-e^{-2\alpha_{p}}}.$ (S.34)
@@
which establishes (S.29). Combining (S.28) and (S.29) proves (S.26), which in turn confirms (S.24).
To prove (S.25), we show that when $\gamma_{p}\gtrsim 1$,
$\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\gtrsim\big(1-\gamma_{p})^{2}.$ (S.35)
When $p(1-\gamma_{p})^{3}$ is sufficiently large, if (S.35) holds, then using the fact that $\big|w_{j}^{k,m_{p},p}-\tilde{w}_{j}^{k,\gamma_{p}}\big|\lesssim\frac{1}{p(1-\gamma_{p})}$, we know
@@
Noting that $\gamma_{p}\gtrsim$ implies $\alpha_{p}\gtrsim 1$, we obtain (S.35).
Therefore, the claim is proven for (S.24) and (S.25). To upper bound the sum $\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}$ from (S.25), we use the following rearranged version of recurrence (S.4):
$1-\tilde{w}_{j}^{k,\gamma_{p}}=e^{-\alpha_{p}k}(1-\tilde{w}_{j-1}^{k,\gamma_{p}})+e^{-\alpha_{p}(k-j+1)}\tilde{w}_{j-1}^{k,\gamma_{p}}.$ (S.36)
Upon squaring both sides of (S.36), we obtain
@@
$\qquad 2e^{-\alpha_{p}(2k-j+1)}\tilde{w}_{j-1}^{k,\gamma_{p}}(1-\tilde{w}_{j-1}^{k,\gamma_{p}})+e^{-2\alpha_{p}(k-j+1)}\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}.$ (S.37)
Summing (S.1.3) over $1\leq j\leq k$, using the Cauchy-Schwarz inequality, and the fact from Theorem 3.1 that $\tilde{w}_{j-1}^{k,\gamma_{p}}\geq\tilde{w}_{j}^{k,\gamma_{p}}$, we obtain
@@
$\qquad+2e^{-\alpha_{p}k}\sqrt{\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}}$ (S.38)
@@
Solving the quadratic inequality (S.1.3) yields
$\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\leq\frac{\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}}{(1-e^{-\alpha_{p}k})^{2}}.$ (S.39)
Now, from the upper bound (10), we have $\big(\tilde{w}_{j-1}^{k,\gamma_{p}}\big)^{2}\leq(1-e^{-\alpha_{p}k})^{2}$ for all $j\geq 1$. Substituting this into (S.39), we obtain the bound
$\sum_{j=1}^{k}\big(1-\tilde{w}_{j}^{k,\gamma_{p}}\big)^{2}\leq\sum_{j=1}^{k}e^{-2\alpha_{p}(k-j+1)}=\frac{e^{-2\alpha_{p}}(1-e^{-2\alpha_{p}k})}{1-e^{-2\alpha_{p}}}.$ (S.40)
@@
$A=\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}],\qquad B=\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}^{2}_{j,p}],$ (S.41)
and note that $A\geq B$ by applying Definition 3.1. Using (S.41) and by the previous inequalities in (S.40) and (S.34), the training error gap (S.21) is at least
@@
$\qquad\geq\frac{3}{2}(1-\epsilon)Be^{-\alpha_{p}}\frac{1-e^{-\alpha_{0}}}{1+e^{-\alpha_{0}}}-(1+\epsilon)Ae^{-2\alpha_{p}},$ (S.42)